Find the shortest route from node 1 to node 7 in the network shown. If the constant is 1 it must be entered in the box. IfOptimal Solution: Variable Value X12 X13 X14 X23 X25 X32 X35 X46 X52 X53 X56 X57 X65 X67 Shortest Route: Length =
Answer #1

Answer:

Min = 11,12 +9:13 +19:14 +9:23 +11,25 +9:32 + 5.35 +18,46 +11,52 +5.53 + 20:56 +17:57 +20,65 +1767

using excel solver

SUM X fo =SUM(C12:112) D E B C F G H I LM 8 100 100 100 100 100 100 100 100 100 100 100 100 100 100 200 100 100 100 Total aua

Some of the values entered are dummy values. Values must be any between 100 or more.

SUM - X fx =SUM(C12:C18) D E B C F G H I J K L M 3 5 100 2 11 100 4 19 100 100 7 100 100 100 100 100 100 18 100 100 100 1 100

SUM X fo=J12-C19 D E B C F G H I J K L M 23 4 5 9 100 100 100 100 100 i 100 Total 6 7 Total Out Out - In 0=312-014! 0 0 Total

SUM A X L = SUMPRODUCT(C3:19,C12:118) D E F G B C H I J K L M N 3 4 5 7 1 100 2 119 6 100 100 100 9 100 100 9 100 100 100 100

Me Oyun Of To! O by Charynge $C$125B 100 100 le: 100 100 100 Sinct to the Congress 11 300 9 100 100 100 100 100 1005 100 100

B C H 100 100 100 20 100 100 Total 6 7 Total Out Out - In 0 = 0 -1 = = 0 O 00 oo 1 Total In 1 0 1

Variable Value x12 0 x13 x14 x23 X25 x32 X35 x46 x52 x53 x56 x57 x65 X67

The shortest path is 1-3-5-7 and it takes 31 to complete